OpenStudy (anonymous):

Two boats start their journey from the same point A and travel along directions AC and AD, as shown below. What is the distance, CD, between the boats?

OpenStudy (anonymous):

Do you know how to solve 30-60-90 triangles? The sides are in this ratio: short side = 1, hypotenuse = 2, third side = sqrt(3)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So you can tell me what BC is?

OpenStudy (anonymous):

\[\tan(60) = \frac{100}{BC}\] Find BC from here..

OpenStudy (anonymous):

BC=57.7 right?

OpenStudy (anonymous):

In that triangle, BC is to 100 as 1 is to sqrt(3)

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You can do it other way also: \[\tan(60) = \sqrt{3}\]

OpenStudy (anonymous):

@Sunshine447 do you know about sines and cosines?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

For 30-60-90 I prefer not to get into that, just remember the basic ratio 1-2-sqrt(3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[BC = \frac{100}{\sqrt{3}} \implies \frac{100 \sqrt{3}}{3}\] Now find: \[\tan(30) = \frac{100}{BD}\]

OpenStudy (anonymous):

\[\tan(30) = \frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

Now, you can get BD because that triangle is also 30-60-90

OpenStudy (anonymous):

Then just do: \(BD - BC\) You will get \(CD\)..

OpenStudy (anonymous):

115.5?

OpenStudy (anonymous):

Let me check..

OpenStudy (anonymous):

sqrt(3) 100 - 100/sqrt(3)

OpenStudy (anonymous):

I got >>> a = 3**0.5 >>> 1.0/a 0.5773502691896258 >>> a*100 - 1.0/a 172.6277304876981

OpenStudy (anonymous):

\[BD = \frac{300}{\sqrt{3}} \implies \frac{300 \sqrt{3}}{3}\]

OpenStudy (anonymous):

oops >>> a*100 - 100.0/a 115.47005383792515 >>> we match

OpenStudy (anonymous):

yay! thanks guys!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[BD - BC = \frac{200 \sqrt{3}}{3} \implies CD = 115.48 \approx \color{green}{115.5}\]